# Bounds on the charge of the graviton using gravitational wave observations

### v1

@@ I Introduction
The long-range character of gravity demands that the quanta of gravity, gravitons, must be massless. To phrase it differently, in general relativity (GR), gravitational waves (GWs) travel at the speed of light, asserting that gravitons must have zero rest mass. However, in search of new physics, we must aim to test this key idea and look for potential deviations. The detection of GWs from merging compact binary systems Abbott et al. (2016a, 2019a, 2021a, 2023) by the LIGO-Virgo-KAGRA (LVK) collaboration Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021) has opened up a new avenue for testing such essential foundations of Einstein’s theory of general relativity Abbott et al. (2016b, 2019b, 2019c, 2021b, 2021c).
@@
To investigate any potential deviations from GR, one could consider specific alternative models or theories and constrain their parameters from observations using the gravitational waveform calculated within the theory; such an analysis will allow specific interpretations of the results within the framework of that theory. However, calculating waveforms in alternative theories of gravity has proved to be difficult. For this reason, one resorts to performing theory-agnostic studies where deviations from the predictions of GR are quantified through phenomenological parameters. For example, suppose we consider GWs generated in an alternative theory of gravity that lacks a mathematical framework to comprehend the GW generation process. Nevertheless, since the GWs propagate over cosmological length scales, cumulative corrections affecting the GW propagation could become dominant over other effects related to its generation. Such propagation effects offer an excellent possibility for theory-agnostic tests of new gravitational physics and allow for stringent bounds on parameters quantifying the violation of general relativity. A specific illustration of such an approach involves constraining the mass of the graviton through the GW observations, as presented in Will (1998); Mirshekari et al. (2012); Abbott et al. (2016b, 2021b, 2021c). To constrain the graviton mass, one computes the time difference of detection, $\Delta t_{o}$, of two gravitons emitted at a time difference $\Delta t_{e}$ apart at the source due to an additional mass-dependent term in the dispersion relation of the GWs Will (1998). This non-zero mass adds an additional contribution to the GW phasing at $-1$ post-Newtonian (PN) order. Using this modified phasing formula, the latest GW observations bound the mass of the graviton ($m_{g}$) to be $m_{g}\leq 1.27\times 10^{-23}\,{\rm eV}/c^{2}$ Abbott et al. (2021c).
@@
In a similar phenomenological spirit, let us ask whether the graviton can be endowed with a non-zero electric charge. Here, we note earlier attempts at constraining the electric charge associated with neutrinos and photons Barbiellini and Cocconi (1987); Raffelt (1994); Altschul (2007, 2008) using astrophysical observations. These observations utilized the interaction of the particle with the ambient magnetic field and the resulting additional time delay for their bounds. Interestingly, the relative time delay induced by the mass and the electric charge of a particle of energy $E$ has the same $E^{-2}$ scaling Raffelt (1994); Barbiellini and Cocconi (1987); Cocconi (1988). This suggests that if we consider a massive graviton, the measured mass may be indistinguishable from the possible electric charge of the graviton.
@@
It should be noted that attributing an electric $U(1)$ charge to a massless particle is not straightforward. The mass has a purely classical interpretation in terms of the dispersion relation of the propagating wave, whereas charge is inherently associated with the quanta of the field. The electric charge of the graviton will be associated with a coupling of the gravitational field to the background electromagnetic field via an $U(1)$ interaction. Attributing such a charge to a massless graviton may lead to inconsistencies. For instance, attempts at constructing quantum theories of higher spin fields coupled to the $U(1)$ gauge field resulted in inconsistencies in the quantization of such theories if the fields are massless Berends et al. (1979); de Wit and Freedman (1980); Johnson and Sudarshan (1961). Additionally, the Weinberg-Witten theorem argues that in a higher spin ($j>1/2$; $j=2$ for graviton) theory that allows the construction of a Lorentz-covariant conserved four current, one cannot have a charged massless particle Weinberg and Witten (1980). However, there are no apparent field theoretic obstructions on a massive graviton like the one considered in Will (1998); Abbott et al. (2016b, 2021b, 2021c) having a $U(1)$ charge.
@@
The above observation, in conjunction with the possible degeneracy of the effect of putative mass with the electric charge of the graviton, makes the phenomenology of the graviton charge particularly relevant. Similar to a charged neutrino or photon, a charged massive graviton will interact with the intergalactic magnetic field, possibly altering its propagation. In this work, we shall consider the effects of an electrically charged massive graviton on the GW phasing and impose simultaneous bounds on its charge and mass. We shall compute the energy-dependent time delay induced on the propagation of a charged massive graviton due to its interaction with the intergalactic magnetic field in the same spirit as the effect of a non-zero charge on the propagation of the neutrino Barbiellini and Cocconi (1987) or the photon Raffelt (1994); Altschul (2007). This time delay can be associated with the observed GW phasing through the methods developed in Will (1998); here, we explicitly demonstrate the degeneracy of the graviton charge and mass. Following this, we use the LVK observations to impose simultaneous constraints on the graviton’s mass and charge. Our analysis results in an upper bound on the graviton charge in terms of the electron charge $e$ of $|q_{g}|/e<3\times 10^{-34}{}$. Furthermore, based on previous work Altschul (2007, 2008), we show that, in the presence of a charged graviton, GWs detected at two detectors separated by a finite distance would have a phase difference; this phase difference can be understood as the Aharonov-Bohm phase Aharonov and Bohm (1959) due to the interaction of the charged particle with the intergalactic magnetic field. Using the measurement of the phase difference (consistent with zero) between the LIGO Hanford and LIGO Livingston detectors from GW190814, we place an independent bound of $|q_{g}|/e<2\times 10^{-26}{}$ on the charge of the graviton.
II Imprint of charge on the waveform
@@
Let us consider compact binaries with component masses $m_{1}$ and $m_{2}$, which radiate away energy through GWs, where we assume the radius of their circular orbit to vary adiabatically. This problem is well understood perturbatively within GR using the framework of restricted post-Newtonian (PN) formalism, which expresses the amplitude $A(f)=\mathcal{A}\,f^{-7/6}$ through the quadrupole approximation and the phase $\psi(f)$ is given by an expansion in powers of ${\beta(f):=v(f)/c}$ Blanchet et al. (2002); Blanchet (2014):
$h(f)=\mathcal{A}\,f^{-7/6}\,e^{i\,\psi(f)}\,,$ (1)
@@
We shall follow the prescription outlined in Ref. Will (1998) for our discussion. We will consider gravitational radiation emitted from such a binary with a frequency $f_{e}$ and the graviton associated with this radiation. If the graviton has a non-zero charge, its trajectory will be curved when exposed to a transverse magnetic field. This alteration of the trajectory will affect the time the graviton takes to reach the detector once emitted. This delay will be in addition to the delay caused by a non-zero mass, considered in Will (1998).
To compute the additional time delay caused by an alteration of the trajectory, we will perform an analysis similar to the neutrino charge estimation Barbiellini and Cocconi (1987). Since the correlation length associated with the intergalactic magnetic field (IGM) is much shorter than the distance of propagation of the graviton, we need to account for the presence of multiple magnetic field domains along the path of propagation Altschul (2007, 2008).
@@
To start, we note that if the IGM has a typical correlation length of $\lambda_{B}$, the region between the source and the observer can be considered to be divided into $N$ magnetic field domains, each of length $\lambda_{B}$ such that the luminosity distance, $D_{L}=N\lambda_{B}$. A magnetic field of magnitude $(B)$ transverse to the direction of propagation will result in the trajectory of a charged graviton being curved, with a radius of curvature given by $r=\frac{p}{Bq_{g}}$, $p$ being the momentum and $q_{g}$ being the possible charge of the graviton. This has been illustrated in Fig. (1). The additional time in relation to the uncharged case that a charged massive graviton would take due to the presence of a transverse magnetic field to travel the distance $\lambda_{B}$ can now be calculated.
By referring the schematic in Fig. (1), the additional time a graviton would take to traverse the curved path from $P1$ to $P2$ of length $L=r\theta$, $\theta$ being the angle subtended by the radii is given to be of the following form when $\theta$ is small Barbiellini and Cocconi (1987)
$\delta t_{\lambda}$ $=\frac{(L-\lambda_{B})}{v_{g}}$ (2)
$\approx\frac{\lambda_{B}^{2}}{4!\,r^{2}}\Delta t_{\lambda}\leavevmode\nobreak\ .$
@@
$\Delta t_{\lambda}$ $=\frac{\lambda_{B}}{v_{g}}=\frac{\lambda_{B}E_{g}}{pc^{2}}$ (3)
$E_{g}=hf_{e}$ is the energy of the graviton with frequency $f_{e}$. Using the above expression for $\Delta t_{\lambda}$, Eq. (2) can be written as
$\delta t_{\lambda}$ $=\frac{\lambda_{B}^{3}}{4!\,cr^{2}\sqrt{1-\frac{m_{g}^{2}c^{4}}{(hf_{e})^{2}}}}=\frac{c\,\lambda_{B}^{3}B^{2}q_{g}^{2}}{4!\,h^{2}f_{e}^{2}\left(1-\frac{m_{g}^{2}c^{4}}{h^{2}f_{e}^{2}}\right)^{\frac{3}{2}}},$ (4)
@@
$\delta t_{q}\approx\sum_{N}p_{\lambda}\delta t_{\lambda}\leavevmode\nobreak\ .$ (5)
Where $p_{\lambda}$ is the probability for the magnetic field inside the magnetic field domain to be aligned transverse to the direction of propagation, and for randomly oriented magnetic field domains this can be approximated as the probability for it to be aligned along four of the six cardinal directions, so $p_{\lambda}\approx\frac{2}{3}$. Now, we can use Eq. (4) to write Eq. (5) as
$\delta t_{q}$ $\approx\frac{2N}{3}\frac{c\lambda_{B}^{3}B^{2}q_{g}^{2}}{4!h^{2}f_{e}^{2}(1-\frac{m_{g}^{2}c^{4}}{h^{2}f_{e}^{2}})^{\frac{3}{2}}}\approx\frac{2}{3}\frac{c\,D_{L}\lambda_{B}^{2}B^{2}q_{g}^{2}}{4!\,h^{2}f_{e}^{2}\left(1-\frac{m_{g}^{2}c^{4}}{h^{2}f_{e}^{2}}\right)^{\frac{3}{2}}}\leavevmode\nobreak\ .$ (6)
Having computed the additional time delay [^1] We note that there might be an $\mathcal{O}(1)$ factor correction to the above expression that arises as a consequence of the assumptions made, like the IGM being of a fixed magnitude and suddenly changing direction at the edge of each magnetic field domain and equating $p_{\lambda}$ to $2/3$ Altschul (2008, 2007). However, such corrections will not change the order of magnitude bound on the graviton charge. a charged massive graviton would encounter while traversing a distance $D_{L}$, we will next compute the phase correction, which is expected at a GW detector Will (1998); Ghosh et al. (2023).
@@
$\Delta t_{o}$ $\approx\Delta t_{e}+\delta t^{m}_{f_{e},f_{e^{\prime}}}+\delta t^{q}_{f_{e},f_{e^{\prime}}}\leavevmode\nobreak\ .$ (7)
Where, $\delta t^{m}_{f_{e},f_{e^{\prime}}}$, is the expected time difference between two gravitons emitted at orbital frequencies $f_{e}$ and $f_{e^{\prime}}$, as a result of the graviton being massive without any charge Will (1998).
$\delta t^{m}_{f_{e},f_{e^{\prime}}}$ $\approx\frac{D_{L}c^{3}m_{g}^{2}}{2h^{2}}\left(\frac{1}{f_{e}^{2}}\right)-\frac{D_{L}c^{3}m_{g}^{2}}{2h^{2}}\left(\frac{1}{f_{e^{\prime}}^{2}}\right).$ (8)
In the above expression, we used the assumption $m_{g}c^{2}<<hf_{e}$. While $\delta t^{q}_{f_{e},f_{e^{\prime}}}$, is the correction due to the possible change in the path of a charged and massive graviton. Using Eq. (6), we can compute $\delta t^{q}_{f_{e},f_{e^{\prime}}}$ to be
$\delta t^{q}_{f_{e},f_{e^{\prime}}}$ $\approx\frac{2}{3}\frac{cD_{L}\lambda_{B}^{2}B^{2}q_{g}^{2}}{4!h^{2}}\left(\frac{1}{f_{e}^{2}}\right)-\frac{2}{3}\frac{cD_{L}\lambda_{B}^{2}B^{2}q_{g}^{2}}{4!h^{2}}\left(\frac{1}{f_{e^{\prime}}^{2}}\right)\leavevmode\nobreak\ .$ (9)
After assuming that the evolution of the binary is driven by gravitational back-reaction such that the associate $df_{e}/dt_{e}$ is well approximated by general relativity Will (1998); Ghosh et al. (2023) and any correction to the generation of gravitational radiation due to the modified theory of gravity is negligible in relation to the terms proportional to the large luminosity distance $D_{L}$ we will get the following equation (see Appendix A) for the GW phase at the detector.
$\psi(f)=$ $2\pi f\tilde{t}_{c}-\tilde{\Phi}_{c}-\frac{\pi}{4}+\sum_{n}\alpha^{n}f^{\frac{2n}{3}}$ (10)
 $\text{\hskip 42.67912pt}-\left\{\frac{\pi m_{g}^{2}c^{3}}{h^{2}}+\left(\frac{2}{3}B^{2}\lambda_{B}^{2}\right)\frac{2\pi q_{g}^{2}c}{4!h^{2}}\right\}\frac{D_{L}}{f},$
@@
Comparing the GW phasing given in Eq. (10) for a charged massive graviton with that of a massive graviton in Eqn. 3.8 of Will (1998) suggests that if we entertain the possibility of $U(1)$ charge on the graviton, the estimates on the mass of the graviton $m_{g}$ can be interpreted as a bound on the effective mass, $m_{\rm eff}$ of a charged massive graviton given by
$m_{\rm eff}=\sqrt{m_{g}^{2}+\left(\frac{2}{3}B^{2}\lambda_{B}^{2}\right)\frac{2q_{g}^{2}}{4!c^{2}}}\leavevmode\nobreak\ .$ (11)
One may argue strong magnetic fields can exist in the vicinity of a binary merger and those need to be taken into account while calculating the GW phasing in the presence of a charged graviton. However, even if these magnetic fields are strong they would not be expected to stay coherent over long length scales like the intergalactic magnetic field, thus minimizing their effect on the GW phasing. Hence, we will assume that the $B$ and $\lambda_{B}$ that enters Eq. (10) are the same as their values in the intergalactic magnetic field.
III Detector dependent phase shifts due to graviton propagation
In addition to the above de-phasing of the GW resulting from a non-zero graviton charge, we can also expect a relative frequency-independent phase shift between GW signals from the same source observed by two separate detectors. This is a consequence of the well-known Aharonov-Bohm effect Aharonov and Bohm (1959). A similar analysis has been carried out to bound the charge of a photon Altschul (2007, 2008); here we extend the analysis to bound the graviton charge.

### v3

# S. Nair    A. Vijaykumar    S. Sarkar

@@ 1 Introduction
In general relativity (GR), gravitational waves (GWs) travel at the speed of light, asserting that gravitons must have zero rest mass. However, in search of new physics, we must test this key idea and look for potential deviations. The detection of GWs from merging compact binary systems [1, 2, 3, 4] by the LIGO-Virgo-KAGRA (LVK) collaboration [5, 6, 7] has opened up a new avenue for testing such essential foundations of Einstein’s theory of general relativity [8, 9, 10, 11, 12].
@@
To investigate any potential deviations from GR, one could consider specific alternative models or theories and constrain their parameters from observations using the gravitational waveform calculated within the theory. However, calculating waveforms in alternative theories of gravity has proved to be difficult. But suppose we consider GWs generated in a modified theory of gravity that lacks a mathematical framework to comprehend the GW generation process. Nevertheless, since the GWs propagate over cosmological length scales, cumulative corrections affecting the GW propagation could become dominant over other effects related to its generation. Such propagation effects offer an excellent possibility for theory-agnostic tests of new gravitational physics and allow for stringent bounds on parameters quantifying the violation of general relativity. A specific illustration of such an approach involves constraining the mass of the graviton through the GW observations, as presented in [13, 14, 8, 11, 12]. To constrain the graviton mass, one computes the time difference of detection, $\Delta t_{o}$, of two gravitons emitted at a time difference $\Delta t_{e}$ apart at the source with an additional mass-dependent term in the dispersion relation of the GWs [13]. This non-zero mass adds an additional contribution to the GW phasing at $-1$ post-Newtonian (PN) order. Using this modified phasing formula, the latest GW observations bound the mass of the graviton ($m_{g}$) to be $m_{g}\leq 1.27\times 10^{-23}\,{\rm eV}/c^{2}$ [12].
@@
In a similar phenomenological spirit, let us ask whether the graviton can be endowed with a non-zero electric charge. Here, we note earlier attempts at constraining the electric charge associated with neutrinos and photons [15, 16, 17, 18] using astrophysical observations. These observations utilized the interaction of the particle with the ambient magnetic field and the resulting additional time delay for their bounds. Interestingly, the relative time delay induced by the mass and the electric charge of a particle of energy $E$ has the same $E^{-2}$ scaling [16, 15, 19]. This suggests that if we consider a massive graviton, the measured mass may be indistinguishable from the possible electric charge of the graviton.
@@
It should be noted that attributing an electric $U(1)$ charge to a massless particle is not straightforward. The electric charge of the graviton will be associated with a coupling of the gravitational field to the background electromagnetic field via an $U(1)$ interaction. Attributing such a charge to a massless graviton may lead to inconsistencies. For instance, attempts at constructing quantum theories of higher spin fields coupled to the $U(1)$ gauge field resulted in inconsistencies in the quantization of such theories if the fields are massless [20, 21, 22]. Additionally, the Weinberg-Witten theorem argues that in a higher spin ($j>1/2$; $j=2$ for graviton) theory that allows the construction of a Lorentz-covariant conserved four current, one cannot have a charged massless particle [23]. However, there are no apparent field theoretic obstructions on a massive graviton like the one considered in [13, 8, 11, 12] having a $U(1)$ charge.
@@
The above observation, in conjunction with the possible degeneracy of the effect of putative mass with the electric charge of the graviton, makes the phenomenology of the graviton charge particularly relevant. In this work, we shall consider the effects of an electrically charged massive graviton on the GW phasing and impose simultaneous bounds on its charge and mass. We shall compute the energy-dependent time delay induced on the propagation of a charged massive graviton due to its interaction with the intergalactic magnetic field in the same spirit as the effect of a non-zero charge on the propagation of the neutrino [15] or the photon [16, 17]. This time delay can be associated with the observed GW phasing through the methods developed in [13]; here, we explicitly demonstrate the degeneracy of the graviton charge and mass. Following this, we use the LVK observations to impose simultaneous constraints on the graviton’s mass and charge. Our analysis results in an upper bound on the graviton charge in terms of the electron charge $e$ of $|q_{g}|/e<3\times 10^{-34}{}$. Furthermore, based on previous work [17, 18], we show that, in the presence of a charged graviton, GWs detected at two detectors separated by a finite distance would have a phase difference; this phase difference can be understood as the Aharonov-Bohm phase [24] due to the interaction of the charged particle with the intergalactic magnetic field. Using the measurement of the phase difference (consistent with zero) between the LIGO Hanford and LIGO Livingston detectors from GW190814, we place an independent bound of $|q_{g}|/e<2\times 10^{-26}{}$ on the charge of the graviton.
2 Imprint of charge on the waveform
@@
Let us consider compact binaries with component masses $m_{1}$ and $m_{2}$, which radiate away energy through GWs, where we assume the radius of their circular orbit to vary adiabatically. This problem is well understood perturbatively within GR using the framework of restricted post-Newtonian (PN) formalism, which expresses the amplitude $A(f)=\mathcal{A}\,f^{-7/6}$ through the quadrupole approximation and the phase $\psi(f)$ is given by an expansion in powers of ${\beta(f):=v(f)/c}$ [25, 26]:
$h(f)=\mathcal{A}\,f^{-7/6}\,e^{i\,\psi(f)}\,,$ (2.1)
@@
We shall follow the prescription outlined in Ref. [13] for our discussion. We will consider gravitational radiation emitted from such a binary with a frequency $f_{e}$ and the graviton associated with this radiation. If the graviton has a non-zero charge, its trajectory will be curved when exposed to a transverse magnetic field. This alteration of the trajectory will affect the time the graviton takes to reach the detector once emitted. This delay will be in addition to the delay caused by a non-zero mass, considered in [13].
To compute the additional time delay caused by an alteration of the trajectory, we will perform an analysis similar to the neutrino charge estimation [15]. Since the correlation length associated with the intergalactic magnetic field (IGM) is much shorter than the distance of propagation of the graviton, we need to account for the presence of multiple magnetic field domains along the path of propagation [17, 18].
@@
To start, we note that if the intergalactic medium has a magnetic field of magnitude $B$ and correlation length $\lambda_{B}$, the region between the source and the observer can be considered to be divided into $N$ magnetic field domains, each of length $\lambda_{B}$ such that the luminosity distance, $D_{L}=N\lambda_{B}$. A magnetic field of magnitude $(B)$ transverse to the direction of propagation will result in the trajectory of a charged graviton being curved, with a radius of curvature given by $r=\frac{p}{Bq_{g}}$, $p$ being the momentum and $q_{g}$ being the possible charge of the graviton. This has been illustrated in Fig. (1). The additional time in relation to the uncharged case that a charged massive graviton would take due to the presence of a transverse magnetic field to travel the distance $\lambda_{B}$ can now be calculated.
By referring the schematic in Fig. (1), the additional time a graviton would take to traverse the curved path from $P1$ to $P2$ of length $L=r\theta$, $\theta$ being the angle subtended by the radii is given to be of the following form when $\theta$ is small [15]
$\delta t_{\lambda}$ $=\frac{(L-\lambda_{B})}{v_{g}}$ (2.2)
$\approx\frac{\lambda_{B}^{2}}{4!\,r^{2}}\Delta t_{\lambda}~{}.$
@@
$\Delta t_{\lambda}$ $=\frac{\lambda_{B}}{v_{g}}=\frac{\lambda_{B}E_{g}}{pc^{2}}$ (2.3)
$E_{g}=hf_{e}$ is the energy of the graviton with frequency $f_{e}$. Using the above expression for $\Delta t_{\lambda}$, Eq. (2.2) can be written as
$\delta t_{\lambda}$ $=\frac{\lambda_{B}^{3}}{4!\,cr^{2}\sqrt{1-\frac{m_{g}^{2}c^{4}}{(hf_{e})^{2}}}}=\frac{c\,\lambda_{B}^{3}B^{2}q_{g}^{2}}{4!\,h^{2}f_{e}^{2}\left(1-\frac{m_{g}^{2}c^{4}}{h^{2}f_{e}^{2}}\right)^{\frac{3}{2}}},$ (2.4)
@@
$\delta t_{q}\approx\sum_{N}p_{\lambda}\delta t_{\lambda}~{}.$ (2.5)
Where $p_{\lambda}$ is the probability for the magnetic field inside the magnetic field domain to be aligned transverse to the direction of propagation, and for randomly oriented magnetic field domains this can be approximated as the probability for it to be aligned along four of the six cardinal directions, so $p_{\lambda}\approx\frac{2}{3}$. Now, we can use Eq. (2.4) to write Eq. (2.5) as
$\delta t_{q}$ $\approx\frac{2N}{3}\frac{c\lambda_{B}^{3}B^{2}q_{g}^{2}}{4!h^{2}f_{e}^{2}(1-\frac{m_{g}^{2}c^{4}}{h^{2}f_{e}^{2}})^{\frac{3}{2}}}\approx\frac{2}{3}\frac{c\,D_{L}\lambda_{B}^{2}B^{2}q_{g}^{2}}{4!\,h^{2}f_{e}^{2}\left(1-\frac{m_{g}^{2}c^{4}}{h^{2}f_{e}^{2}}\right)^{\frac{3}{2}}}~{}.$ (2.6)
Having computed the additional time delay [^1] We note that there might be an $\mathcal{O}(1)$ factor correction to the above expression that arises as a consequence of the assumptions made, like the IGM being of a fixed magnitude and suddenly changing direction at the edge of each magnetic field domain and equating $p_{\lambda}$ to $2/3$ [18, 17]. However, such corrections will not change the order of magnitude bound on the graviton charge. a charged massive graviton would encounter while traversing a distance $D_{L}$, we will next compute the phase correction, which is expected at a GW detector [13, 27].
@@
$\Delta t_{o}$ $\approx\Delta t_{e}+\delta t^{m}_{f_{e},f_{e^{\prime}}}+\delta t^{q}_{f_{e},f_{e^{\prime}}}~{}.$ (2.7)
Where, $\delta t^{m}_{f_{e},f_{e^{\prime}}}$, is the expected time difference between two gravitons emitted at orbital frequencies $f_{e}$ and $f_{e^{\prime}}$, as a result of the graviton being massive without any charge [13].
$\delta t^{m}_{f_{e},f_{e^{\prime}}}$ $\approx\frac{D_{L}c^{3}m_{g}^{2}}{2h^{2}}\left(\frac{1}{f_{e}^{2}}\right)-\frac{D_{L}c^{3}m_{g}^{2}}{2h^{2}}\left(\frac{1}{f_{e^{\prime}}^{2}}\right).$ (2.8)
In the above expression, we used the assumption $m_{g}c^{2}<<hf_{e}$. While $\delta t^{q}_{f_{e},f_{e^{\prime}}}$, is the correction due to the possible change in the path of a charged and massive graviton. Using Eq. (2.6), we can compute $\delta t^{q}_{f_{e},f_{e^{\prime}}}$ to be
$\delta t^{q}_{f_{e},f_{e^{\prime}}}$ $\approx\frac{2}{3}\frac{cD_{L}\lambda_{B}^{2}B^{2}q_{g}^{2}}{4!h^{2}}\left(\frac{1}{f_{e}^{2}}\right)-\frac{2}{3}\frac{cD_{L}\lambda_{B}^{2}B^{2}q_{g}^{2}}{4!h^{2}}\left(\frac{1}{f_{e^{\prime}}^{2}}\right)~{}.$ (2.9)
After assuming that the evolution of the binary is driven by gravitational back-reaction such that the associate $df_{e}/dt_{e}$ is well approximated by general relativity [13, 27] and any correction to the generation of gravitational radiation due to the modified theory of gravity is negligible in relation to the terms proportional to the large luminosity distance $D_{L}$ we will get the following equation (see Appendix A) for the GW phase at the detector.
$\psi(f)=$ $2\pi f\tilde{t}_{c}-\tilde{\Phi}_{c}-\frac{\pi}{4}+\sum_{n}\alpha^{n}f^{\frac{2n}{3}}-\left\{\frac{\pi m_{g}^{2}c^{3}}{h^{2}}+\left(\frac{2}{3}B^{2}\lambda_{B}^{2}\right)\frac{2\pi q_{g}^{2}c}{4!h^{2}}\right\}\frac{D_{L}}{f},$ (2.10)
@@
Comparing the GW phasing given in Eq. (2.10) for a charged massive graviton with that of a massive graviton in Eqn. 3.8 of [13] suggests that if we entertain the possibility of $U(1)$ charge on the graviton, the estimates on the mass of the graviton $m_{g}$ can be interpreted as a bound on the effective mass, $m_{\rm eff}$ of a charged massive graviton given by
$m_{\rm eff}=\sqrt{m_{g}^{2}+\left(\frac{2}{3}B^{2}\lambda_{B}^{2}\right)\frac{2q_{g}^{2}}{4!c^{2}}}~{}.$ (2.11)
One may argue strong magnetic fields can exist in the vicinity of a binary merger and those need to be taken into account while calculating the GW phasing in the presence of a charged graviton. However, even if these magnetic fields are strong they would not be expected to stay coherent over long length scales like the intergalactic magnetic field, thus minimizing their effect on the GW phasing. Hence, we will assume that the $B$ and $\lambda_{B}$ that enters Eq. (2.10) are the same as their values in the intergalactic magnetic field.
3 Detector dependent phase shifts due to graviton propagation
In addition to the above de-phasing of the GW resulting from a non-zero graviton charge, we can also expect a relative frequency-independent phase shift between GW signals from the same source observed by two separate detectors. This is a consequence of the well-known Aharonov-Bohm effect [24]. A similar analysis has been carried out to bound the charge of a photon [17, 18]; here we extend the analysis to bound the graviton charge.